# Weak coloring numbers of minor-closed graph classes

### v1

@@ 5. The base case: Graphs with no $\mathcal{F}$-rich model of a given tree
    $\sigma$ extends $\sigma_{j}$, for every $j\in\{0,\dots,\ell\}$;
@@
Note the similarity of this ordering to the one described in Figure 10. For convenience, let $P_{j}=R_{j}^{\prime}=\emptyset$ for every integer $j$ with $j>\ell$.
Let $u\in V(G)$. We will show that $|\operatorname{WReach}_{r}[G,S,\sigma,u]|\leqslant 5(\delta(k,d+1)+3)\cdot r\log r$. Let $j_{u}\in\{0,\dots,\ell\}$ be such that if $u\in S$, then $u\in P_{j_{u}}$, and otherwise, $N_{G}(V(C))\subseteq P_{j_{u}}\cup P_{j_{u}+1}$, where $C$ is the component of $u$ in $G-S$ (such $j_{u}$ exists by 30.(c)).
@@
Let $\alpha=\max\{a\in I_{s}\mid a\leqslant j_{u}\}$. This is well-defined since $0\in I_{s}$. Let $\beta=\alpha+2^{s}$. Thus if $\beta\leqslant\ell$ then $\beta\in I_{s}$. Next, we argue that
@@
Suppose to the contrary that there is a vertex $v\in\operatorname{WReach}_{r}[G,S,\sigma,u]\cap S_{s}$ with $v\not\in R^{\prime}_{\alpha}\cup R^{\prime}_{\beta}$ and $v\in R^{\prime}_{\gamma}$ for some $\gamma\in I_{s}-\{\alpha,\beta\}$. Then either $\gamma<\alpha$ or $\gamma>\beta$. First, assume $\gamma<\alpha$. Since $R^{\prime}_{\gamma}\subseteq\bigcup_{a\in\{0,\dots,\gamma\}}P_{a}$, by ($\star$)5, every $u$-$v$ path in $G$ intersects $P_{a}$ for each $a\in\{\gamma,\dots,j_{u}\}$. Thus, we have $\mathrm{dist}_{G}(u,v)\geqslant j_{u}-\gamma\geqslant\alpha-\gamma\geqslant 2^{s}>r$, which contradicts the fact that $v\in\operatorname{WReach}_{r}[G,S,\sigma,u]$. Finally, assume that $\gamma>\beta$. If $\beta>\ell$, then the assertion is clear, thus assume $\beta\leqslant\ell$. Since $R^{\prime}_{\gamma}\subseteq\bigcup_{a\in\{\gamma-2^{s}+1,\dots,\gamma\}}P_{a}$, by ($\star$)5, every $u$-$v$ path in $G$ intersects $P_{\beta}$. However, for every $w\in P_{\beta}$, we have $w<_{\sigma}v$, thus, $v\notin\operatorname{WReach}_{r}[G,S,\sigma,u]$, which is a contradiction. We obtain that $\operatorname{WReach}_{r}[G,S,\sigma,u]\cap S_{s}\subseteq R^{\prime}_{\alpha}\cup R^{\prime}_{\beta}$.
For every $\delta\in I_{s}$, by definition of $\sigma$, we have
$\operatorname{WReach}_{r}[G,S,\sigma,u]\cap R^{\prime}_{\delta}\subseteq\operatorname{WReach}_{r}[G-\bigcup_{a\in\{0,\dots,j_{u}-2^{s}\}\cap I_{s}}R^{\prime}_{a},R^{\prime}_{\delta},\sigma_{\delta},u]$
@@
$|\operatorname{WReach}_{r}[G,S,\sigma,u]\cap R^{\prime}_{\delta}|\leqslant(\delta(k,d+1)+3)\cdot r.$
@@
Let $\alpha=\max\{a\in I_{i+1}\mid a\leqslant j_{u}\}$ and $\beta=\alpha+2^{i+1}$. Let $C$ be the component of $u$ in $G-S_{i+1}$. Since $P_{\alpha},P_{\beta}\subseteq S_{i+1}$, by ($\star$)5, $V(C)\cap S\subseteq\bigcup_{a\in\{\alpha+1,\dots,\beta-1\}}P_{a}$. We deduce that
$\operatorname{WReach}_{r}[G-S_{i+1},S-S_{i+1},\sigma|_{S-S_{i+1}},u]\cap(S_{i}-S_{i+1})\subseteq\bigcup_{a\in\{\alpha+1,\dots,\beta-1\}}P_{a}.$
Since the only multiple of $2^{i}$ in $\{\alpha+1,\dots,\beta-1\}$ is $\gamma=\alpha+2^{i}$, we in fact have
$\operatorname{WReach}_{r}[G-S_{i+1},S-S_{i+1},\sigma|_{S-S_{i+1}},u]\cap(S_{i}-S_{i+1})\subseteq R_{\gamma}^{\prime}.$
If $\gamma>\ell$, then $R_{\gamma}^{\prime}=\emptyset$ and if $\gamma\leqslant\ell$, then
$|\operatorname{WReach}_{r}[G-S_{i+1},S-S_{i+1},\sigma|_{S-S_{i+1}},u]\cap(S_{i}-S_{i+1})|$ $\leqslant\operatorname{wcol}_{r}(G-S_{i+1},R^{\prime}_{\gamma},\sigma_{\gamma})$
@@
This concludes the claim that $|\operatorname{WReach}_{r}[G,S,\sigma,u]\cap(S_{i}-S_{i+1})|\leqslant(\delta(k,d+1)+3)\cdot r$.
Since $S=S_{0}$, it follows that setting $S_{s+1}=\emptyset$, we have
@@
Let $k,h,d$ be positive integers with $h\geqslant 2$. There is an integer $c_{0}(h,d,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{h,d}$, then there is a set $S\subseteq V(G)$ such that
@@
Let $G$ be a connected $K_{k}$-minor-free graph and let $\mathcal{F}$ be a family of nonnull connected subgraphs of $G$. Suppose that $G$ has no $\mathcal{F}$-rich model of $F_{h,d}$. Let $r$ be an integer with $r\geqslant 2$. Let $\mathcal{F}^{\prime}$ be the family of all the connected subgraphs $H$ of $G$ such that $H$ contains an $\mathcal{F}|_{H}$-rich model of $F_{h-1,d+1}$. We claim that there is no $\mathcal{F}^{\prime}$-rich model of $F_{2,d}$ in $G$. Suppose to the contrary that $\big{(}B_{x}\mid x\in V(F_{2,d})\big{)}$ is such a model. Let $s$ be the root of $F_{2,d}$ and let $s^{\prime}$ be the root of $F_{h-1,d}$. For every $x\in V(F_{2,d})-\{s\}$, by Lemma 32, there is an $\mathcal{F}$-rich model $\big{(}C_{y}\mid y\in V(F_{h-1,d})\big{)}$ of $F_{h-1,d}$ in $G[B_{x}]$ such that $C_{s^{\prime}}$ contains a vertex of $N_{G}(B_{s})\cap B_{x}$. The union of these models together with $B_{s}$ yields an $\mathcal{F}$-rich model of $F_{h,d}$ in $G$, which is a contradiction. See Figure 11.
@@ 6. Proof of the main theorem
Let $k$ and $t$ be positive integers with $t\geqslant 2$. Let $X$ be a graph with $\operatorname{rtd}_{2}(X)\leqslant t$. There exists an integer $c(t,X,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $X$, then there exists $S\subseteq V(G)$ such that
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$. There is an integer $c_{1}(t,Y,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every $u\in V(G)$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $K_{1}\oplus Y$, then there exist $S\subseteq V(G)$, a tree $T$ rooted in $s\in V(T)$, and a tree partition $\big{(}T,(P_{x}\mid x\in V(T))\big{)}$ of $G[S]$ with $P_{s}=\{u\}$ such that
@@
    for every $x\in V(T)$ let $T_{x}$ be the subtree of $T$ rooted in $x$, then
    $\operatorname{wcol}_{r}\left(G\left[\bigcup_{y\in V(T_{x})}P_{y}\right],P_{x}\right)\leqslant c_{1}(t,X,k)\cdot r^{t-2}\log r$
    for every integer $r$ with $r\geqslant 2$.
@@
Proof of the claim. Let $X^{\prime}=K_{1}\sqcup Y$ and let $c_{1}(t,Y,k)=c(t-1,X^{\prime},k)+3$. We proceed by induction on $|V(G)|$. If $|V(G)|=1$, then since $c_{1}(t,Y,k)\geqslant 1$, the result holds. Next, suppose that $|V(G)|>1$.
@@
    for every $x\in V(T_{C})$ let $T_{C,x}$ be the subtree of $T_{C}$ rooted in $x$, then
    $\operatorname{wcol}_{r}\left(G_{C}\left[\bigcup_{y\in V(T_{C,x})}P_{C,y}\right],P_{C,x}\right)\leqslant c_{1}(t,X,k)\cdot r^{t-2}\log r$
    for every integer $r$ with $r\geqslant 2$.
Then let $S=\bigcup_{C\in\mathcal{C}}S_{C}$, let $T$ be the tree obtained from the disjoint union of all the $T_{C}$ for $C\in\mathcal{C}$ by identifying all the vertices in $\{s_{C}\mid C\in\mathcal{C}\}$ into a single vertex $s$. Finally, let $P_{s}=\{u\}$ and $P_{x}=P_{C,x}$ for every $C\in\mathcal{C}$ and $x\in V(T_{C})-\{s_{C}\}$. Then we claim that (a)-(d) hold. Indeed, for every $F\in\mathcal{F}$, $F$ is connected so either $F$ is a subgraph of $G-\{u\}$ or $F$ contains $u$. In both cases, we see that $V(F)\cap S\neq\emptyset$, so (a) holds. Item (b) holds as $u\in S_{C}$ and $G[S_{C}]$ is connected for all $C\in\mathcal{C}$. Item (c) holds as every component of $G-S$ is a component of $G_{C}-S_{C}$ for some $C\in\mathcal{C}$. Finally, for all $x\in V(T)$ with $x\neq s$ item (d) follows directly from (d’) and the construction of $T$, while when $x=s$, $\operatorname{wcol}_{r}\left(G\left[\bigcup_{y\in V(T)}P_{y}\right],P_{s}\right)=1$ for every integer $r$ with $r\geqslant 2$. From now on we assume that $G-\{u\}$ is connected.
Let $\mathcal{F}^{\prime}$ be the family of all the nonnull connected subgraphs $H$ of $G-\{u\}$ such that $u\in N_{G}(V(H))$ and $F\subseteq H$ for some $F\in\mathcal{F}$. We argue that there is no $\mathcal{F}^{\prime}$-rich model of $X^{\prime}$ in $G-\{u\}$. Indeed, such a model would contain an $\mathcal{F}$-rich model of $K_{1}\sqcup Y$ in $G-\{u\}$ such that every branch set is adjacent to $u$. By adding $u$ to the branch set corresponding to $K_{1}$ in $K_{1}\sqcup Y$, we obtain an $\mathcal{F}$-rich model of $K_{1}\oplus Y$ in $G$, which is a contradiction. This proves that there is no $\mathcal{F}^{\prime}$-rich model of $X^{\prime}$ in $G-\{u\}$.
@@
    for every $x\in V(T_{C})$ let $T_{C,x}$ be the subtree of $T_{C}$ rooted in $x$, then
    $\operatorname{wcol}_{r}\left(G_{C}\left[\bigcup_{y\in V(T_{C,x})}P_{C,y}\right],P_{C,x}\right)\leqslant c_{1}(t,X,k)\cdot r^{t-2}\log r$
    for every integer $r$ with $r\geqslant 2$.
@@
Finally, we argue (d). Let $r$ be an integer with $r\geqslant 2$ and let $x\in V(T)$. For $x=s$, $|P_{s}|=1$, thus the assertion is clear. For $x=s^{\prime}$, we have $T_{s^{\prime}}=V(G)-\{u\}$. By (C’), $\operatorname{wcol}_{r}(G-\{u\},S_{0})\leqslant c(t-1,X^{\prime},k)\cdot r^{t-2}\log r$. Since $Q$ is a geodesic in $G$, by Observation 15,
@@
For $x\in V(T_{C}-\{s_{C}\})$ for some $C\in\mathcal{C}_{1}$, we have $T_{x}=T_{C,x}$, thus, the asserted inequality follows from (d”). This ends the proof of the claim. $\lozenge$
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$. There is an integer $c_{2}(t,Y,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $K_{1}\oplus Y$, then there exists $S\subseteq V(G)$ such that
@@
Proof of the claim. Let $c_{2}(t,Y,k)=2\cdot c_{1}(t,Y,k)$ and let $r$ be an integer with $r\geqslant 2$. We apply 1 with an arbitrary vertex $u\in V(G)$ to obtain $S\subseteq V(G)$, a tree $T$ rooted in $s\in V(T)$, and a tree partition $\big{(}T,(P_{x}\mid x\in V(T))\big{)}$ of $S$ in $G$ with $P_{s}=\{u\}$ such that
@@
    for every $x\in V(T)$ let $T_{x}$ be the subtree of $T$ rooted in $x$, then
    $\operatorname{wcol}_{r}\left(G\left[\bigcup_{y\in V(T_{x})}P_{y}\right],P_{x}\right)\leqslant c_{1}(t,X,k)\cdot r^{t-2}\log r$
    for every integer $r$ with $r\geqslant 2$.
@@
Let $y\in A$. Since $\sigma$ extends $\sigma_{y}$ and $x\in T_{y}$, by 1.(d), we have
$|\operatorname{WReach}_{r}[G,S,\sigma,u]\cap P_{y}|\leqslant\operatorname{wcol}_{r}\left(G\left[\bigcup_{z\in V(T_{y})}P_{z}\right],P_{y}\right)\leqslant c_{1}(t,X,k)\cdot r^{t-2}\log r.$
@@
$\leqslant(r+1)\cdot c_{1}(t,X,k)\cdot r^{t-2}\log r$
$\leqslant 2r\cdot c_{1}(t,X,k)\cdot r^{t-2}\log r=c_{2}(t,Y,k)\cdot r^{t-1}\log r.$
This concludes the proof of (c), and the claim follows. $\lozenge$
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$ and let $h,d$ be positive integers. There exists an integer $c_{3}(Y,h,d,k)$ such that for every connected $K_{k}$-minor-free graph $G$ and for every family $\mathcal{F}$ of nonnull connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $T^{\prime}_{h,d}(Y)$, then there exists $S\subseteq V(G)$ such that
@@
Proof of the claim. We proceed by induction on $h$. When $h=1$, $T^{\prime}_{h,d}(Y)=K_{1}\oplus(Y\sqcup Y)$ and the result follows from the previous claim applied to $Y\sqcup Y$ (note that $\operatorname{rtd}_{2}(Y\sqcup Y)=\operatorname{rtd}_{2}(Y)\leqslant t-1$) by setting $c_{3}(Y,1,d,k)=c_{2}(t,Y\sqcup Y,k)$.
@@
    $\operatorname{wcol}_{r}(G,S_{0})\leqslant c_{2}(t,1,Z,k)\cdot r^{t-1}\log r$ for every integer $r$ with $r\geqslant 2$.

### v2

@@ 5. The base case: Graphs with no $\mathcal{F}$-rich model of a given tree
    $\sigma$ extends $\sigma_{j}$, for every $j\in\{1,\dots,\ell\}$;
@@
Note the similarity of this ordering to the one described in Figure 10. For convenience, let $R^{\prime}_{0}=\emptyset$ and $W_{j}=R^{\prime}_{j}=\emptyset$ for every integer $j$ with $j>\ell$.
We now show (c). Let $u\in V(G)$. We will show that $|\operatorname{WReach}_{r}[G,S,\sigma,u]|\leqslant 5(\delta(k,d+1)+3)\cdot r\log r$. Let $j_{u}\in\{0,\dots,\ell\}$ be minimum such that $u\in W_{j_{u}}$.
@@
Let $\alpha=\max\{0\}\cup\{a\in I_{s}\mid a\leqslant j_{u}\}$, and let $\beta=\alpha+2^{s}$. Thus, if $\beta\leqslant\ell$, then $\beta\in I_{s}$. Next, we argue that
@@
Suppose to the contrary that there is a vertex $v\in\operatorname{WReach}_{r}[G,S,\sigma,u]\cap S_{s}$ with $v\not\in R^{\prime}_{\alpha}\cup R^{\prime}_{\beta}$. Let $\gamma\in I_{s}-\{\alpha,\beta\}$ be such that $v\in R^{\prime}_{\gamma}$. Then either $\gamma<\alpha$, or $\gamma>\beta$. First assume that $\gamma<\alpha$. Since $R^{\prime}_{\gamma}\subseteq\bigcup_{a\in\{0,\dots,\gamma-1\}}W_{a}$ and because $(W_{0},\dots,W_{\ell})$ is a path decomposition of $G$, every $u$-$v$ path in $G$ intersects $W_{a-1}\cap W_{a}$ for each $a\in\{\gamma,\dots,j_{u}\}$. Since $(W_{a-1}\cap W_{a})_{a\in\{1,\dots,\ell\}}$ are pairwise disjoint, we deduce that $\mathrm{dist}_{G}(u,v)\geqslant j_{u}-\gamma\geqslant\alpha-\gamma\geqslant 2^{s}>r$, which contradicts the fact that $v\in\operatorname{WReach}_{r}[G,S,\sigma,u]$. Finally, assume $\gamma>\beta$. Note that $\gamma\leqslant\ell$ since $R^{\prime}_{\gamma}\neq\emptyset$ as $v\in R^{\prime}_{\gamma}$. Since $R^{\prime}_{\gamma}\subseteq\bigcup_{a\in\{\gamma-2^{s},\dots,\gamma-1\}}W_{a}$, and because $(W_{0},\dots,W_{\ell})$ is a path decomposition of $G$, every $u$-$v$ path in $G$ intersects $W_{\beta-1}\cap W_{\beta}$. However, for every $w\in W_{\beta-1}\cap W_{\beta}$, we have $w<_{\sigma}v$, thus, $v\not\in\operatorname{WReach}_{r}[G,S,\sigma,u]$, which is a contradiction. We obtain that $\operatorname{WReach}_{r}[G,S,\sigma,u]\cap S_{s}\subseteq R^{\prime}_{\alpha}\cup R^{\prime}_{\beta}$.
For every $\varepsilon\in\{\alpha,\beta\}$, by definition of $\sigma$, we have
$\operatorname{WReach}_{r}[G,S,\sigma,u]\cap R^{\prime}_{\varepsilon}$ $\subseteq\operatorname{WReach}_{r}\left[G-\textstyle\bigcup_{a\in I_{s}\cap\{1,\dots,\alpha-1\}}R^{\prime}_{a},R^{\prime}_{\varepsilon},\sigma_{\varepsilon},u\right]$
$\subseteq\operatorname{WReach}_{r}\left[G-\textstyle\bigcup_{a\in\{1,\dots,j_{u}-2^{s}\}\cap I_{s}}R^{\prime}_{a},R^{\prime}_{\varepsilon},\sigma_{\varepsilon},u\right]$
@@
$|\operatorname{WReach}_{r}[G,S,\sigma,u]\cap R^{\prime}_{\varepsilon}|\leqslant(\delta(k,d+1)+3)\cdot r.$
@@
Let $\alpha=\max\{a\in I_{i+1}\mid a\leqslant j_{u}\}$ and $\beta=\alpha+2^{i+1}$. Let $C$ be the component of $u$ in $G-S_{i+1}$. Since $W_{\alpha-1}\cap W_{\alpha},W_{\beta-1}\cap W_{\beta}\subseteq S_{i+1}$, and because $(W_{0},\dots,W_{\ell})$ is a path decomposition of $G$, $V(C)\cap S\subseteq\bigcup_{a\in\{\alpha,\dots,\beta-1\}}W_{a}$. We deduce that
$\operatorname{WReach}_{r}[G-S_{i+1},S-S_{i+1},\sigma|_{S-S_{i+1}},u]\cap(S_{i}-S_{i+1})\subseteq\bigcup_{a\in\{\alpha,\dots,\beta-1\}}W_{a}.$
Since the only members of $I_{i}-I_{i+1}$ in $\{\alpha+1,\dots,\beta-1\}$ is $\gamma=\alpha+2^{i}$, we in fact have
$\operatorname{WReach}_{r}[G-S_{i+1},S-S_{i+1},\sigma|_{S-S_{i+1}},u]\cap(S_{i}-S_{i+1})\subseteq R^{\prime}_{\gamma},$
and we deduce that
$|\operatorname{WReach}_{r}[G-S_{i+1},S-S_{i+1},\sigma|_{S-S_{i+1}},u]\cap(S_{i}-S_{i+1})|$ $\leqslant\operatorname{wcol}_{r}(G-S_{i+1},R_{\gamma}^{\prime},\sigma_{\gamma})$
@@
For convenience let $S_{s+1}=\emptyset$. Since $S=S_{0}$, it follows that
@@
Let $k,h,d$ be positive integers with $h\geqslant 2$. There is an integer $c_{0}(h,d,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $F_{h,d}$, then there is a set $S\subseteq V(G)$ such that
@@
Let $G$ be a connected $K_{k}$-minor-free graph and let $\mathcal{F}$ be a family of connected subgraphs of $G$. Suppose that $G$ has no $\mathcal{F}$-rich model of $F_{h,d}$. Let $r$ be an integer with $r\geqslant 2$. Let $\mathcal{F}^{\prime}$ be the family of all the connected subgraphs $H$ of $G$ such that $H$ contains an $\mathcal{F}|_{H}$-rich model of $F_{h-1,d+1}$. We claim that there is no $\mathcal{F}^{\prime}$-rich model of $F_{2,d}$ in $G$. Suppose to the contrary that $\big{(}B_{x}\mid x\in V(F_{2,d})\big{)}$ is such a model. Let $s$ be the root of $F_{2,d}$ and let $s^{\prime}$ be the root of $F_{h-1,d}$. For every $x\in V(F_{2,d})-\{s\}$, by Lemma 32, there is an $\mathcal{F}$-rich model $\big{(}C_{y}\mid y\in V(F_{h-1,d})\big{)}$ of $F_{h-1,d}$ in $G[B_{x}]$ such that $C_{s^{\prime}}$ contains a vertex of $N_{G}(B_{s})\cap B_{x}$. The union of these models together with $B_{s}$ yields an $\mathcal{F}$-rich model of $F_{h,d}$ in $G$, which is a contradiction. See Figure 11.
@@ 6. Proof of the main theorem
Let $k$ and $t$ be positive integers with $t\geqslant 2$. Let $X$ be a graph with $\operatorname{rtd}_{2}(X)\leqslant t$. There exists an integer $c(t,X,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $X$, then there exists $S\subseteq V(G)$ such that
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$. There is an integer $c_{1}(t,Y,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every $u\in V(G)$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $K_{1}\oplus Y$, then there exist $S\subseteq V(G)$, a tree $T$ rooted in $s\in V(T)$, and a tree partition $\big{(}T,(P_{x}\mid x\in V(T))\big{)}$ of $G[S]$ with $P_{s}=\{u\}$ such that
@@
    for every $x\in V(T)$,
    $\operatorname{wcol}_{r}\left(G_{x},P_{x}\right)\leqslant c_{1}(t,Y,k)\cdot r^{t-2}\log r$
    for every integer $r$ with $r\geqslant 2$, where, for $T_{x}$ being the subtree of $T$ rooted in $x$, $G_{x}$ is the subgraph of $G$ induced by the union of $U_{x}=\bigcup_{y\in V(T_{x})}P_{y}$ and all the vertex sets of the components of $G-S$ having a neighbor in $U_{x}$.
@@
Proof of the claim.
Let $X^{\prime}=K_{1}\sqcup Y$ and let $c_{1}(t,Y,k)=c(t-1,X^{\prime},k)+3$. We proceed by induction on $|V(G)|$. If $|V(G)|=1$, then since $c_{1}(t,Y,k)\geqslant 1$, the result holds. Next, suppose that $|V(G)|>1$.
@@
    for every $x\in V(T_{C})$,
    $\operatorname{wcol}_{r}\left(G_{C,x},P_{C,x}\right)\leqslant c_{1}(t,Y,k)\cdot r^{t-2}\log r$
    for every integer $r$ with $r\geqslant 2$, where, for $T_{C,x}$ being the subtree of $T_{C}$ rooted in $x$, $G_{C,x}$ is the subgraph of $G_{C}$ induced by the union of $U_{C,x}=\bigcup_{y\in V(T_{C,x})}P_{C,y}$ and all the vertex sets of the components of $G_{C}-S_{C}$ having a neighbor in $U_{C,x}$
Then let $S=\bigcup_{C\in\mathcal{C}}S_{C}$, let $T$ be the tree obtained from the disjoint union of all the $T_{C}$ for $C\in\mathcal{C}$ by identifying all the vertices in $\{s_{C}\mid C\in\mathcal{C}\}$ into a single vertex $s$. Finally, let $P_{s}=\{u\}$ and $P_{x}=P_{C,x}$ for every $C\in\mathcal{C}$ and $x\in V(T_{C})-\{s_{C}\}$. Then we claim that (a)-(d) hold. Indeed, for every $F\in\mathcal{F}$, $F$ is connected so either $F$ is a subgraph of $G-\{u\}$ or $F$ contains $u$. In both cases, we see that $V(F)\cap S\neq\emptyset$, so (a) holds. Item (b) holds as $u\in S_{C}$ and $G[S_{C}]$ is connected for all $C\in\mathcal{C}$. Item (c) holds as every component of $G-S$ is a component of $G_{C}-S_{C}$ for some $C\in\mathcal{C}$. Finally, for all $x\in V(T)$ with $x\neq s$ item (d) follows directly from (d’) and the construction of $T$, while when $x=s$, $\operatorname{wcol}_{r}\left(G_{s},P_{s}\right)=1$ for every integer $r$ with $r\geqslant 2$. From now on we assume that $G-\{u\}$ is connected.
Let $\mathcal{F}^{\prime}$ be the family of all the connected subgraphs $H$ of $G-\{u\}$ such that $u\in N_{G}(V(H))$ and $F\subseteq H$ for some $F\in\mathcal{F}$. We argue that there is no $\mathcal{F}^{\prime}$-rich model of $X^{\prime}$ in $G-\{u\}$. Suppose by contradiction that it exists. Such a model would contain an $\mathcal{F}$-rich model of $K_{1}\sqcup Y$ in $G-\{u\}$ such that every branch set is adjacent to $u$. By adding $u$ to the branch set corresponding to $K_{1}$ in $K_{1}\sqcup Y$, we obtain an $\mathcal{F}$-rich model of $K_{1}\oplus Y$ in $G$, which is a contradiction. This proves that there is no $\mathcal{F}^{\prime}$-rich model of $X^{\prime}$ in $G-\{u\}$.
@@
    for every $x\in V(T_{C})$,
    $\operatorname{wcol}_{r}\left(G_{C,x},P_{C,x}\right)\leqslant c_{1}(t,Y,k)\cdot r^{t-2}\log r$
    for every integer $r$ with $r\geqslant 2$, where, for $T_{C,x}$ being the subtree of $T_{C}$ rooted in $x$, $G_{x}$ is the subgraph of $G$ induced by the union of $U_{x}=\bigcup_{y\in V(T_{x})}P_{y}$ and all the vertex sets of the components of $G_{C}-S_{C}$ having a neighbor in $U_{x}$.
@@
Finally, we argue (d). For every $x\in V(T)$, we denote by $T_{x}$ the subtree of $T$ rooted in $x$, and by $U_{x}$ the subgraph of $G$ induced by $U_{x}=\bigcup_{y\in V(T_{x})}P_{y}$ with the vertex sets of all the components of $G-S$ having a neighbor in $U_{x}$. Let $r$ be an integer with $r\geqslant 2$ and let $x\in V(T)$. For $x=s$, $|P_{s}|=1$, thus the assertion is clear. For $x=s^{\prime}$, we have $G_{s^{\prime}}$ is a union of components of $G-\{u\}$. By (C’), $\operatorname{wcol}_{r}(G_{s^{\prime}},S_{0})=\operatorname{wcol}_{r}(G-\{u\},S_{0})\leqslant c(t-1,X^{\prime},k)\cdot r^{t-2}\log r$. Since $Q$ is a geodesic in $G$, by Observation 15,
@@
For $x\in V(T_{C}-\{s_{C}\})$ for some $C\in\mathcal{C}_{1}$, we have $T_{x}=T_{C,x}$, thus, the asserted inequality follows from (d”). This ends the proof of the claim. ∎
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$. There is an integer $c_{2}(t,Y,k)$ such that for every connected $K_{k}$-minor-free graph $G$, for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $K_{1}\oplus Y$, then there exists $S\subseteq V(G)$ such that
@@
Proof of the claim.
Let $c_{2}(t,Y,k)=2\cdot c_{1}(t,Y,k)$ and let $r$ be an integer with $r\geqslant 2$. We apply 1 with an arbitrary vertex $u\in V(G)$ to obtain $S\subseteq V(G)$, a tree $T$ rooted in $s\in V(T)$, and a tree partition $\big{(}T,(P_{x}\mid x\in V(T))\big{)}$ of $S$ in $G$ with $P_{s}=\{u\}$ such that
@@
    for every $x\in V(T)$,
    $\operatorname{wcol}_{r}\left(G_{x},P_{x}\right)\leqslant c_{1}(t,Y,k)\cdot r^{t-2}\log r$
    for every integer $r$ with $r\geqslant 2$, where, for $T_{x}$ being the subtree of $T$ rooted in $x$, the graph $G_{x}$ is the union of $U_{x}=\bigcup_{y\in V(T_{x})}P_{y}$ with the vertex sets of all the components of $G-S$ having a neighbor in $U_{x}$.
@@
Let $y\in A$. Since $\sigma$ extends $\sigma_{y}$ and $x_{u}\in V(T_{y})$, by 1.(d), we have
$|\operatorname{WReach}_{r}[G,S,\sigma,u]\cap P_{y}|\leqslant\operatorname{wcol}_{r}\left(G_{y},P_{y}\right)\leqslant c_{1}(t,Y,k)\cdot r^{t-2}\log r.$
@@
$\leqslant(r+1)\cdot c_{1}(t,Y,k)\cdot r^{t-2}\log r$
$\leqslant 2r\cdot c_{1}(t,Y,k)\cdot r^{t-2}\log r=c_{2}(t,Y,k)\cdot r^{t-1}\log r.$
This concludes the proof of (c), and the claim follows. ∎
@@
Let $Y$ be a graph with $\operatorname{rtd}_{2}(Y)\leqslant t-1$ and let $h,d$ be positive integers. There exists an integer $c_{3}(Y,h,d,k)$ such that for every connected $K_{k}$-minor-free graph $G$ and for every family $\mathcal{F}$ of connected subgraphs of $G$, if $G$ has no $\mathcal{F}$-rich model of $T^{\prime}_{h,d}(Y)$ [^9] See the definition on page 4., then there exists $S\subseteq V(G)$ such that
@@
Proof of the claim.
We proceed by induction on $h$. When $h=1$, $T^{\prime}_{h,d}(Y)=K_{1}\oplus(Y\sqcup Y)$ and the result follows from the previous claim applied to $Y\sqcup Y$ (note that $\operatorname{rtd}_{2}(Y\sqcup Y)=\operatorname{rtd}_{2}(Y)\leqslant t-1$) by setting $c_{3}(Y,1,d,k)=c_{2}(t,Y\sqcup Y,k)$.
@@
    $\operatorname{wcol}_{r}(G,S_{0})\leqslant c_{2}(t,Z,k)\cdot r^{t-1}\log r$ for every integer $r$ with $r\geqslant 2$.